OpenStudy (anonymous):

Solve for the indicated variable. A=h/2 (a+b); for h

OpenStudy (anonymous):

A=h/2(a+b) Multiply both sides by 2(a+b) We get 2(a+b)A= h Now multiply out 2(a+b)A We get 2Aa +2Ab =h Divide both sides by A We get 2Aa+ 2Ab = h Now take 2Ab to the right side 2Aa = h -2Ab a= (h -2Ab)/2A

OpenStudy (anonymous):

that's incorrect @xxccjxx

OpenStudy (anonymous):

Do you mean how do you solve for the variable a? That would just indicate we need to isolate a. So first, multiply both sides by 2/h to move it to the left. 2A / h = a+b subtract b a=(2A / h) - b